OpenStudy (anonymous):

Solve: 9x2 – 60x = -100

OpenStudy (anonymous):

10/3

OpenStudy (anonymous):

yes its 10/3 or is it 10/3 and -10/3

OpenStudy (anonymous):

9x^2 -60x+100=0 is your equation , so if we factorize we'll get (3x-10)^2=0 so x=10/3

OpenStudy (campbell_st):

well rewrite the equation as \[9x^2 - 60x + 100 = 0\] this equation is a perfect square \[(ax - b)^2 = a^x^2 - 2abx + b^2\] you just need to find a and b when you get them just solve ax - b = 0 for the solutions.